OpenStudy (anonymous):

completely factor the following expression; 100y^2 -36z^2

hartnn (hartnn):

note : \(10^2=100 \quad and \quad6^2 = 36\) hint: use \(a^2-b^2=(a+b)(a-b)\) does this help ??

OpenStudy (anonymous):

okay i got 4(25y^2-9z^2)

OpenStudy (anonymous):

but i think i need to factor 25y^2-9z^2 also but im not sure if it factors anymore

hartnn (hartnn):

yes,it does. thats step is also the correct first step. now note 25=5^2 and 9=3^2 so its (5y)^2-(3z)^2 now can u factorize using a^2-b^2 ?

OpenStudy (anonymous):

4(25y^2-9z^2) and then factor 25y^2-9z^2 = (5y^2)(2)-(3z^2)(3) is this right

hartnn (hartnn):

where does (3) come from ?? in the end and u still need to factorize we know \(a^2-b^2=(a+b)(a-b)\) so just put a=5y^2 and b= 3z up there and tell me what u get.

hartnn (hartnn):

*a=5y

OpenStudy (anonymous):

a^2-b^2=(5y+3z)(5y-3z)

hartnn (hartnn):

are u getting what i wanted to say?

OpenStudy (anonymous):

yes is what i just type right

hartnn (hartnn):

yup thats correct :) and don't forget 4 in the end 4(5y+3z)(5y-3z)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes thank you

hartnn (hartnn):

welcome :)